OpenStudy (anonymous):

Medal How many kilograms of nickel silver containing 16% zinc and how many kilograms of nickel silver containing 31% zinc must be melted together to produce 711 kg of a new nickel silver containing 24% zinc? Treat the percents given in this exercise as exact numbers, and work to three significant digits. ? kg of 16% alloy and ? kg of 31% alloy

OpenStudy (anonymous):

@campbell_st can you please help me

OpenStudy (anonymous):

we need a variable, lets say \(x\) is the number of kg of 31% zink

OpenStudy (anonymous):

since the total is \(711\) the amount of 21% zink must therefore be \(711-x\)

OpenStudy (anonymous):

16% i meant

OpenStudy (anonymous):

then \(x\) kilos of 31% zink has \(.31x\) zink in it. and similarly the \(711-x\) kilos of 16% zink has \(.16(711-x)\) kilos of zinkk

OpenStudy (anonymous):

you want \(711\times .24=170.64\) kilos of zink set \[.31x+.16(711-x)=170.64\] and solve for \(x\)

OpenStudy (anonymous):

easier to solve by removing annoying decimals and solve\[31x+16(711-x)=17064\]

OpenStudy (anonymous):

31x+11376-16x=17064

OpenStudy (anonymous):

x=379.2

OpenStudy (anonymous):

am I right @satellite73